# Wenhao Wang

### v2

@@ Abstract
This paper explores in-context learning for image copy detection (ICD), i.e., prompting an ICD model to identify replicated images with new tampering patterns without the need for additional training. The prompts (or the contexts) are from a small set of image-replica pairs that reflect the new patterns and are used at inference time. Such in-context ICD has good realistic value, because it requires no fine-tuning and thus facilitates fast reaction against the emergence of unseen patterns. To accommodate the “seen $\rightarrow$ unseen” generalization scenario, we construct the first large-scale pattern dataset named AnyPattern, which has the largest number of tamper patterns ($90$ for training and $10$ for testing) among all the existing ones. We benchmark AnyPattern with popular ICD methods and reveal that existing methods barely generalize to novel patterns. We further propose a simple in-context ICD method named ImageStacker. ImageStacker learns to select the most representative image-replica pairs and employs them as the pattern prompts in a stacking manner (rather than the popular concatenation manner). Experimental results show (1) training with our large-scale dataset substantially benefits pattern generalization ($+26.66\%$ $\mu AP$), (2) the proposed ImageStacker facilitates effective in-context ICD (another round of $+16.75\%$ $\mu AP$), and (3) AnyPattern enables in-context ICD, i.e. without such a large-scale dataset, in-context learning does not emerge even with our ImageStacker. Beyond the ICD task, we also demonstrate how AnyPattern can benefit artists, i.e., the pattern retrieval method trained on AnyPattern can be generalized to identify style mimicry by text-to-image models. The project is publicly available at https://anypattern.github.io.
Figure 1: Top: The comparison between the standard updating process of Image Copy Detection (ICD) and the proposed in-context ICD. Unlike the standard updating approach, our in-context ICD eliminates the need for fine-tuning, making it more efficient. Bottom: AnyPattern is the first large-scale pattern dataset, featuring 90 base and 10 novel patterns. Using 90 base patterns, we generate a training dataset containing 10 million images. Note that each pattern in this paper refers to a class of transformations that are diverse within themselves (see Section A in the Appendix).
@@ 1 Introduction
To further demonstrate the significance of introducing the AnyPattern dataset, we present an additional application using AnyPattern with the proposed pattern retrieval method. The text-to-image model can be used to mimic the style of artwork with little cost, and this threatens the livelihoods and creative rights of artists. To help them protect their work, we treat an artist’s ‘style’ as a ‘pattern’ and generalize the trained pattern retrieval method to identify generated images with style mimicry. We provide a detailed discussion of this application of AnyPattern in the Appendix (Section B).
@@
    To further highlight the generalization and importance of AnyPattern, we present another application of AnyPattern, i.e. the pattern retrieval method trained on AnyPattern can be generalized to identify generated images by text-to-image models that most closely matches the style of a given real artwork.
@@ 2 Related Works
In-context Learning for Computer Vision. In-context learning originates from large language models like GPT-3 [1] and Instruction GPT [2]. In the pure computer vision area, in-context learning is a relatively new concept and presents challenges for implementation. The earliest known work, MAE-VQGAN [3], implements in-context learning through image inpainting. Following MAE-VQGAN [3], (Un)SupPR [4] discusses how to find good visual in-context examples; Prompt-SelF [5] explores the factors that affect the performance of visual in-context learning; Painter [6] designs a generalist vision model to automatically generate images according to the example pairs; Prompt Diffusion [7] proposes diffusion-based generative models. Specific computer vision areas also see work such as SegGPT [8] for image segmentation. These methods typically implement in-context learning in a concatenating manner, while we propose a stacking design.
Existing Image Copy Detection Methods. Existing ICD methods [9, 10, 11, 12, 13, 14] can be broadly categorized into contrastive learning-based algorithms and deep metric learning-based algorithms. The use of contrastive learning for training an ICD model is natural due to its reliance on data augmentation. For example, SSCD [9] is based on the InfoNCE [15] loss and introduces a differential entropy regularization to differentiate nearby vectors. CNNCL [10] employs a large memory bank with a contrastive loss to learn from numerous positive and negative pairs. The application of deep metric learning to ICD is intuitive given that ICD is, fundamentally, a retrieval task. EfNet [11] proposes a “drip training” procedure, whereby the number of classes used to train a model is incrementally increased. BoT [12] sets a robust baseline for ICD and introduces descriptor stretching, which normalizes scores at the feature level. However, none of these methods can be directly applied to our in-context ICD scenario as they never consider generalizing their models to novel patterns, let alone using only image-replica pairs.
@@ 3.2 AnyPattern Dataset
The source images to generate these replicated images and prompts are from the DISC21 dataset [16, 17]. DISC21 has $1$ million unlabeled training images, from which we randomly select $100,000$ images as the un-edited images following [12, 13]. Each training image is transformed $99$ times by randomly selected training patterns. Together with the original training images, we construct a training dataset containing $10$ million images. Owing to the complexity and volume of the patterns, this process is distributed across $200$ CPU nodes in a supercomputing cluster and requires about one million CPU core hours. We adopt the gallery dataset from DISC21 as our gallery. The query set includes $25,000$ queries, among which $5,000$ are generated by applying a randomly selected novel pattern combination to gallery images, and the remaining $20,000$ queries serve as distractors (without true matches in the gallery).
@@ 3.3 Comparison against Existing Datasets
Currently, there are three publicly available ICD benchmarks, i.e. CopyDays [18], DISC21 [16, 17], and NDEC [13].
CopyDays [18] was launched in 2009. This dataset contains only 157 query images and 3,000 gallery images, and lacks training data. The types of tampering patterns involved are relatively straightforward, such as alterations in contrast and blurring.
DISC21 [16, 17] was established in 2021 as an extensive benchmark for ICD, notable for its massive scale, including one million training images and one million gallery images, along with complicated tampering patterns. Additionally, it includes numerous distractor queries, which do not correspond to any true matches in the gallery.
NDEC [13] addresses the challenge of hard negatives in ICD, i.e., some images may appear very similar yet are not replications. By incorporating this aspect of hard negatives, NDEC enhances the realism of ICD evaluations.
@@
Beyond the above discussions, we also introduce the application of AnyPattern to identify style mimicry, highlighting its generalizability and importance. Please refer to Section B in the Appendix.
@@ 4.1 Baseline
This section briefly overviews the ICD baseline implemented in our ImageStacker. Following [12, 11], we conceptualize ICD as an image retrieval task, primarily adopting deep metric learning methods. Specifically, we treat each original image and all its replicas as a training class and perform deep metric learning on these classes. Pairwise training [19, 20], classification training [21, 22, 23], or their combination can be utilized for this purpose. In our baseline, we select classification training, specifically CosFace [22], due to its demonstrated effectiveness and simplicity.
@@ 4.2 Pattern Retrieval
Drawing inspiration from image retrieval techniques, we propose a pattern retrieval method for identifying the image-replica pairs corresponding to a given query (Fig. 3 (a)). The training of pattern retrieval can be seen as a multi-label classification task. When using Vision Transformer (ViT) [24] as the pattern extractor, we design a pattern token $x^{0}_{ptr}$ and concatenate it to the ViT input:
@@ 4.3 Stacking
To accommodate these $9$-channel image patches, we initialize a new embedding layer (while keeping the hidden size and all other details of the original ViT). For a query/replica, we use the retrieved image-replica pair or ground truth as the prompt; for a gallery/original image, we directly duplicate itself two times as the pseudo-prompt. The stacked image-replica pair alters the feed-forward process, thus allowing for a conditioned input image feature. The advantage of our stacking design lies in introducing an inductive bias to the in-context learning, which emphasizes the contrasts at the same or similar positions between the original image and the copy. Since many tampering patterns occur at the same or similar positions, this inductive bias brings more benefits to the in-context learning process compared to the traditional concatenation method without any inductive bias. In the Appendix (Section C), we further analyze what ViT learns with our ImageStacker design.
@@ 4.4 Unifying Pattern Retrieval and Stacking
where $\mathcal{L}_{cls}$ is the CosFace loss [22], and $\lambda$ is the balance parameter.
@@
Figure 4: The performance of ICD state-of-the-arts, CLIP [25], and our Basl. (Small) on the DISC21 dataset and novel patterns. Our Basl. (Small) is trained with the same patterns with BoT [12]. Since these algorithms (including our Basl. (Small)) are not designed to handle input of image-replica pairs, we demonstrate their performance decrease by directly testing their trained models on the $10$ novel patterns from AnyPattern.
@@ 5.1 Evaluation Metrics and Training Details
Please refer to the Appendix (Section D and E).
5.2 The Challenge from Novel Patterns
In this section, we present the performance degradation of trained ICD models (SSCD [9], BoT [12], EfNet [11], and CNNCL [10]) and the CLIP [25] model, when encountering novel patterns. All the ICD state-of-the-arts are trained on the DISC21 dataset, and we select the most successful CLIP [25] model, which is trained on the $2$ billion sample English subset of LAION-5B [26] and achieves a zero-shot top-1 accuracy of $80.1\%$ on ImageNet-1k [27]. The corresponding $\mu AP$ and $recall@1$ scores are summarized in Fig. 4, leading us to two main observations. First, all the ICD models, despite being trained with different methods and pattern combinations, experience a significant accuracy decrease (about $60\%$ for $\mu AP$ and $recall@1$) when the evaluation dataset changes from DISC21 to novel patterns. Notably, our baseline is trained on the same pattern combination with BoT [12] (Basl. (Small)) and maintains comparable performance with state-of-the-arts on both the DISC21 dataset and the $10$ novel patterns. Second, while CLIP models display impressive results in zero-shot image classification and image retrieval tasks, they are not ideally suited for ICD tasks (less than $10\%$ $\mu AP$). This can be attributed to CLIP being predominantly trained on natural images. In light of these findings, we argue that in practical scenarios, the continuous emergence of novel patterns poses a significant challenge for deployed ICD models.
Table 1: The performance improvement on the novel patterns from AnyPattern and in-context learning. Please see the performance on base patterns in the Appendix (Section F).
@@ 5.2 The Challenge from Novel Patterns
ImageStacker 57.31 60.14 59.96 57.10 60.86

### v3

@@ Abstract
This paper explores in-context learning for image copy detection (ICD), i.e., prompting an ICD model to identify replicated images with new tampering patterns without the need for additional training. The prompts (or the contexts) are from a small set of image-replica pairs that reflect the new patterns and are used at inference time. Such in-context ICD has good realistic value, because it requires no fine-tuning and thus facilitates fast reaction against the emergence of unseen patterns. To accommodate the “seen $\rightarrow$ unseen” generalization scenario, we construct the first large-scale pattern dataset named AnyPattern, which has the largest number of tamper patterns ($90$ for training and $10$ for testing) among all the existing ones. We benchmark AnyPattern with popular ICD methods and reveal that existing methods barely generalize to novel patterns. We further propose a simple in-context ICD method named ImageStacker. ImageStacker learns to select the most representative image-replica pairs and employs them as the pattern prompts in a stacking manner (rather than the popular concatenation manner). Experimental results show (1) training with our large-scale dataset substantially benefits pattern generalization ($+26.66\%$ $\mu AP$), (2) the proposed ImageStacker facilitates effective in-context ICD (another round of $+16.75\%$ $\mu AP$), and (3) AnyPattern enables in-context ICD, i.e., without such a large-scale dataset, in-context learning does not emerge even with our ImageStacker. Beyond the ICD task, we also demonstrate how AnyPattern can benefit artists, i.e., the pattern retrieval method trained on AnyPattern can be generalized to identify style mimicry by text-to-image models. The project is publicly available at https://anypattern.github.io.
Keywords:
Image Copy Detection AnyPattern In-context Learning Style Mimicry
Figure 1: Top: The comparison between the standard updating process of Image Copy Detection (ICD) and the proposed in-context ICD. Unlike the standard updating approach, our in-context ICD eliminates the need for fine-tuning, making it more efficient. Bottom: AnyPattern is the first large-scale pattern dataset, featuring 90 base and 10 novel patterns. Using 90 base patterns, we generate a training dataset containing 10 million images. Note that each pattern in this paper refers to a class of transformations that are diverse within themselves (see Appendix (Section A)).
@@ 1 Introduction
To further demonstrate the significance of introducing the AnyPattern dataset, we present an additional application using AnyPattern with the proposed pattern retrieval method. The text-to-image diffusion model can be used to mimic the style of artwork with little cost, and this threatens the livelihoods and creative rights of artists. To help them protect their work, we treat an artist’s ‘style’ as a ‘pattern’ and generalize the trained pattern retrieval method to identify generated images with style mimicry.
@@
    To further highlight the generalization and importance of AnyPattern dataset, we present its another application, i.e. the pattern retrieval method trained on AnyPattern can be generalized to identify generated images by text-to-image diffusion models that most closely matches the style of a given real artwork.
@@ 2 Related Works
In-context Learning for Computer Vision. In-context learning originates from large language models like GPT-3 (Brown et al., 2020) and Instruction GPT (Ouyang et al., 2022). In the pure computer vision area, in-context learning is a relatively new concept and presents challenges for implementation. The earliest known work, MAE-VQGAN (Bar et al., 2022), implements in-context learning through image inpainting. Following MAE-VQGAN (Bar et al., 2022), (Un)SupPR (Zhang et al., 2023) discusses how to find good visual in-context examples; Prompt-SelF (Sun et al., 2023) explores the factors that affect the performance of visual in-context learning; Painter (Wang et al., 2023b) designs a generalist vision model to automatically generate images according to the example pairs; Prompt Diffusion (Wang et al., 2023d) proposes diffusion-based generative models. Specific computer vision areas also see work such as SegGPT (Wang et al., 2023c) for image segmentation. These methods typically implement in-context learning in a concatenating manner, while we propose a stacking design.
Existing Image Copy Detection Methods. Existing ICD methods (Pizzi et al., 2022; Yokoo, 2021; Papadakis and Addicam, 2021; Wang et al., 2021, 2023a; Fernandez et al., 2023) can be broadly categorized into contrastive learning-based algorithms and deep metric learning-based algorithms. The use of contrastive learning for training an ICD model is natural due to its reliance on data augmentation. For example, SSCD (Pizzi et al., 2022) is based on the InfoNCE (Oord et al., 2018) loss and introduces a differential entropy regularization to differentiate nearby vectors. CNNCL (Yokoo, 2021) employs a large memory bank with a contrastive loss to learn from numerous positive and negative pairs. The application of deep metric learning to ICD is intuitive given that ICD is, fundamentally, a retrieval task. EfNet (Papadakis and Addicam, 2021) proposes a “drip training” procedure, whereby the number of classes used to train a model is incrementally increased. BoT (Wang et al., 2021) sets a robust baseline for ICD and introduces descriptor stretching, which normalizes scores at the feature level. However, none of these methods can be directly applied to our in-context ICD scenario as they never consider generalizing their models to novel patterns, let alone using only image-replica pairs.
@@ 3.2 AnyPattern Dataset
The source images to generate these replicated images and prompts are from the DISC21 dataset (Douze et al., 2021; Papakipos et al., 2022). DISC21 has $1$ million unlabeled training images, from which we randomly select $100,000$ images as the un-edited images following (Wang et al., 2021, 2023a). Each training image is transformed $99$ times by randomly selected training patterns. Together with the original training images, we construct a training dataset containing $10$ million images. Owing to the complexity and volume of the patterns, this process is distributed across 200 CPU nodes in a supercomputing cluster and requires about one million CPU core hours. We adopt the gallery dataset from DISC21 as our gallery. The query set includes $25,000$ queries, among which $5,000$ are generated by applying a randomly selected novel pattern combination to gallery images, and the remaining $20,000$ queries serve as distractors (without true matches in the gallery).
@@ 3.3 Comparison against Existing Datasets
Currently, there are three publicly available ICD benchmarks, i.e. CopyDays (Douze et al., 2009), DISC21 (Douze et al., 2021), and NDEC (Wang et al., 2023a).
CopyDays (Douze et al., 2009) was launched in 2009. This dataset contains only 157 query images and 3,000 gallery images, and lacks training data. The types of tampering patterns involved are relatively straightforward, such as alterations in contrast and blurring.
DISC21 (Douze et al., 2021; Papakipos et al., 2022) was established in 2021 as an extensive benchmark for ICD, notable for its massive scale, including one million training images and one million gallery images, along with complicated tampering patterns. Additionally, it includes numerous distractor queries, which do not correspond to any true matches in the gallery.
NDEC (Wang et al., 2023a) addresses the challenge of hard negatives in ICD, i.e., some images may appear very similar yet are not replications. By incorporating this aspect of hard negatives, NDEC enhances the realism of ICD evaluations.
@@ 4.1 Baseline
This section briefly overviews the ICD baseline implemented in our ImageStacker. Following (Wang et al., 2021; Papadakis and Addicam, 2021), we conceptualize ICD as an image retrieval task, primarily adopting deep metric learning methods. Specifically, we treat each original image and all its replicas as a training class and perform deep metric learning on these classes. Pairwise training (Sohn, 2016; Hermans et al., 2017), classification training (Liu et al., 2016; Wang et al., 2018; Sun et al., 2020), or their combination can be utilized for this purpose. In our baseline, we select classification training, specifically CosFace (Wang et al., 2018), due to its demonstrated effectiveness and simplicity.
@@ 4.2 Pattern Retrieval
Drawing inspiration from image retrieval techniques, we propose a pattern retrieval method for identifying the image-replica pairs corresponding to a given query (Fig. 3 (a)). The training of pattern retrieval can be seen as a multi-label classification task. When using Vision Transformer (ViT) (Dosovitskiy et al., 2020) as the pattern extractor, we design a pattern token $x^{0}_{ptr}$ and concatenate it to the ViT input:
@@ 4.3 Stacking
To accommodate these $9$-channel image patches, we initialize a new embedding layer (while keeping the hidden size and all other details of the original ViT). For a query/replica, we use the retrieved image-replica pair or ground truth as the prompt; for a gallery/original image, we directly duplicate itself two times as the pseudo-prompt. The stacked image-replica pair alters the feed-forward process, thus allowing for a conditioned input image feature. The advantage of our stacking design lies in introducing an inductive bias to the in-context learning, which emphasizes the contrasts at the same or similar positions between the original image and the copy. Since many tampering patterns occur at the same or similar positions, this inductive bias brings more benefits to the in-context learning process compared to the traditional concatenation method without any inductive bias.
@@ 4.4 Unifying Pattern Retrieval and Stacking
where $\mathcal{L}_{cls}$ is the CosFace loss (Wang et al., 2018), and $\lambda$ is the balance parameter.
@@
5 AnyPattern helps artists
5.1 Background
DreamBooth (Ruiz et al., 2023) presents a novel method for personalizing text-to-image diffusion models. By using a few reference images of a subject, DreamBooth enables the model to generate new, high-quality images of that subject in various contexts specified by textual descriptions. This fine-tuning approach retains the model’s general capabilities while enhancing its ability to produce detailed and contextually appropriate depictions of the specific subject. Due to its potential negative societal impact, with concerns that ‘malicious parties might try to use such images to mislead viewers’, the inventors at Google decided not to release any code or trained models. However, a third party re-implemented it and made it publicly available (Xiao, 2022).
Style mimicry. After the release of DreamBooth, people discover that it can easily be used to mimic the styles of any artist. Specifically, anyone can collect as few as five artworks created by an artist and spend less than one dollar to train a DreamBooth model, which can then generate numerous images in the same style. For instance, Ogbogu Kalu released a tuned DreamBooth model on Hugging Face (ogkalu, 2024), which was tuned using the artworks of six comic artists and has gained significant popularity. In Fig. 4, we contrast the real artworks created by these artists with the images generated by the tuned DreamBooth model, demonstrating that their styles are indeed very similar.
Opinion of artists and others. The phenomenon of style mimicry has sparked a debate about the ethics of fine-tuning AI on the artworks of living artists in the comments on Reddit (user, 2022) and other places. Supporters of AIGC consider the generated images ‘incredibly beautiful’ and describe the released model as their ‘favorite custom model by far’. Meanwhile, some legal professionals argue that it is lawful because ‘style is not copyrightable’. However, there is more to it than that. Firstly, artists feel frustrating, uncomfortable and invasive for this (Baio, 2022). For instance, Hollie Mengert wonders if the model’s creator simply did not think of her as a person. She also has concerns about copyright because some of the training artworks from her are created for copyright holders, such as Disney and Penguin Random House. Secondly, these models may end artists ability to earn a living (Shan et al., 2023). Artists invest significant time and effort in cultivating their unique styles, which is a critical aspect of their livelihood. As a model replicates these styles without offering compensation, artists’ opportunities to market their work and connect with potential buyers are significant hindered. Finally, the artistic creation will be a swan song (Nguyen, 2023). This imitation by AI can be demoralizing for art students who are training to become the next generation of artists. Seeing AI models potentially replace their future roles can be discouraging and impact their career aspirations.
5.2 Our target
Glaze (Shan et al., 2023) offers ‘style cloaks’ to artists to help mislead the mimicry of their styles by text-to-image models. However, the authors of Glaze acknowledge a limitation: this preventative approach can only protect newer artworks. More specifically, many works have already been downloaded from art repositories such as ArtStation and DeviantArt, and these artists’ styles can still be mimicked using older artworks collected before Glaze was released.
Here, we aim to provide artists with such a ‘style retrieval’ tool. It is a direct application of our AnyPattern and pattern retrieval method, showing their generalizability to other datasets or real-world scenarios where patterns significantly differ. Specifically, we treat an artist’s ‘style’ as a ‘pattern’. Therefore, using our pattern retrieval method trained with AnyPattern, we can directly search a database containing millions of generated images to identify the image that most closely matches the style of a given real artwork. A demonstration of such a process is shown in Fig. 5.
5.3 Implementation
Experimental setup. To test the generalizability of our trained pattern retrieval method in identifying style mimicry by text-to-image models, we first construct a database of millions of generated images. Specifically, we utilize prompts from DiffusionDB (Wang et al., 2023e) and Stable Diffusion V1.5 (RunwayML, 2022) to generate $1,819,776$ images. Then, we collect several publicly available artworks by artists such as Charlie Bowater, Hollie Mengert, Mario Alberti, Pepe Larraz, Andreas Rocha, and James Daly III as query images. It is important to note that our use of these artworks falls under the category of fair use for research purposes (contributors, 2024), thereby avoiding any copyright infringement. Employing the style descriptor [^1] https://github.com/WangWenhao0716/AnyPatternStyle we trained on AnyPattern, we extract a 768-dimensional vector from each generated image and real artwork. By computing the cosine similarities, we identify the top-10 generated images that most closely match the style of each real artwork.
5.4 Limitations and future directions
@@
Figure 7: The performance of ICD state-of-the-arts, CLIP (Radford et al., 2021), and our Basl. (Small) on the DISC21 dataset and novel patterns. Our Basl. (Small) is trained with the same patterns with BoT (Wang et al., 2021). Since these algorithms (including our Basl. (Small)) are not designed to handle input of image-replica pairs, we demonstrate their performance decrease by directly testing their trained models on the $10$ novel patterns from AnyPattern.
@@ 6.1 Evaluation Metrics and Training Details
Evaluation metrics. We employ two evaluation metrics for the in-context ICD, namely $\mu AP$ and $recall@1$/$R@1$. $\mu AP$ serves as an overall evaluation metric and is equivalent to the area under the Precision-Recall curve, providing a comprehensive measure of both the precision and recall of our model across varying thresholds. On the other hand, the $recall@1$ metric is query-specific. It checks whether the actual (correct) result appears first in the list of all returned results, offering insight into the effectiveness of our model in accurately retrieving the most relevant result at the top.
Table 1: The performance improvement on the novel patterns from AnyPattern and in-context learning.
@@
ImageStacker $57.31$ $60.14$ $59.96$ $57.10$ $\mathbf{60.86}$
Training details. We implement our ImageStacker using PyTorch (Paszke et al., 2019) and distribute its training across eight Nvidia A100 GPUs. We use ViT-B/16 (Dosovitskiy et al., 2020) as the backbone, which is pre-trained on the ImageNet dataset (Deng et al., 2009) using DeiT (Touvron et al., 2021) unless otherwise specified. Before training, we resize images to a resolution of $224\times 224$ pixels. We set the balance parameter, $\lambda$, at $1$ and use a batch size of $512$. Each batch adopts the standard PK sampling method, with $128$ classes and $4$ images per class. The total number of training epochs is $25$ with a cosine-decreasing learning rate. The margin $m$ and scale $s$ in CosFace loss Wang et al. (2018) are set to $0.35$ and $64$, respectively.
6.2 The Challenge from Novel Patterns
In this section, we present the performance degradation of trained ICD models (SSCD (Pizzi et al., 2022), BoT (Wang et al., 2021), EfNet (Papadakis and Addicam, 2021), and CNNCL (Yokoo, 2021)) and the CLIP (Radford et al., 2021) model, when encountering novel patterns. All the ICD state-of-the-arts are trained on the DISC21 dataset, and we select the most successful CLIP (Radford et al., 2021) model, which is trained on the $2$ billion sample English subset of LAION-5B (Schuhmann et al., 2022) and achieves a zero-shot top-1 accuracy of $80.1\%$ on ImageNet-1k (Deng et al., 2009). The corresponding $\mu AP$ and $recall@1$ scores are summarized in Fig. 7, leading us to two main observations. First, all the ICD models, despite being trained with different methods and pattern combinations, experience a significant accuracy decrease (about $60\%$ for $\mu AP$ and $recall@1$) when the evaluation dataset changes from DISC21 to novel patterns. Notably, our baseline is trained on the same pattern combination with BoT (Wang et al., 2021) (Basl. (Small)) and maintains comparable performance with state-of-the-arts on both the DISC21 dataset and the $10$ novel patterns. Second, while CLIP models display impressive results in zero-shot image classification and image retrieval tasks, they are not ideally suited for ICD tasks (less than $10\%$ $\mu AP$). This can be attributed to CLIP being predominantly trained on natural images. In light of these findings, we argue that in practical scenarios, the continuous emergence of novel patterns poses a significant challenge for deployed ICD models.
[…]
SSCD
BoT
EfNet
CNNCL
Baseline
AnyPattern $77.54$ $76.12$ $77.56$ $79.84$ $79.37$
ImageStacker $81.39$ $79.98$ $80.11$ $84.13$ $83.56$
[…]
SSCD
BoT
EfNet
CNNCL
Baseline
AnyPattern $81.76$ $80.37$ $81.13$ $83.41$ $83.05$
ImageStacker $84.52$ $83.96$ $83.78$ $86.07$ $85.85$